[Brahms-l] Shift report 20040311 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Mar 11 2004 - 15:52:07 EST
------------------------------------------------------------
Date: Thu Mar 11 15:52:06 2004, Shift: 08:00-16:00
Supervisor: BB, with: 
Log book#15 pages 83 - 83

Shift Summary:
------------------------------------------------------------
1. A beam quench occurred 1 hour and 20 minutes into the
   shift. No beam for rest of shift.

2. FS at 8 degrees full field.  MRS at 90 degrees half
   field.
   Run 10904 (begun by previous shifters) Tr2 83k Tr 3 91k
   Run 10906                 Tr 2 32k    Tr3 49k

   There was ample time for overdue pedestal and non TPC 
   pedestal runs.

3. After run 10906 Dana moved the FS to 10 degrees.

   Magnet water circulation was checked by MC. This required
   that the magnets be placed on standby. The FS magnets are
   now back on at 1/4 field and D5 is at 1/2 field.

   A problem with the RunServer was cured with a remedy
   given in the FAQ,or maybe it cured itself of a temporary 
   hiccup.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 11 15:52:44 2004

This archive was generated by hypermail 2.1.8 : Thu Mar 11 2004 - 15:53:09 EST